solve equations let z be the parameter
5x + 3y + 4z=19
3x - y + z = -4

Answer by vleith(2983)

5x + 3y + 4z=19
4z = 19 - 5x - 3y
z = (19 - 5x - 3y)/4

3x - y + z = -4
z = -4 - 3x + y